# Improved High-probability Convergence Guarantees of Decentralized SGD

### v2

@@ 3.3 Strongly Convex Costs
In this subsection we provide improved results for the last iterate of strongly convex costs. It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [nesterov-lectures_on_cvxopt]. Denote the global minimizer by $x^{\star}\in\mathbb{R}^{d}$, noting that $f^{\star}\coloneqq f(x^{\star})$ and let $\kappa\coloneqq\frac{L}{\mu}\geq 1$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively. Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a]. To show similar benefits in the HP sense, a different approach to Lemma 3 is required. This is achieved by carefully bounding the MGF of the consensus gap and leveraging properties of strong convexity. The formal result is stated next.
@@
Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$
@@
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5). Using Lemma 6 from the Appendix, it can be shown that the RHS is of the order $e^{\mathcal{O}(\nu K_{t+1}\alpha_{t}^{2})}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional term in the exponent on the RHS, of order $\mathcal{O}(\nu K_{t+1}\lambda^{t}\Delta_{x})$, which decays geometrically (recall that $\lambda\in[0,1)$). However, we assume same model initialization, for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which allows us to establish sharper bounds and ensure linear speed-up in the number of users is achieved.
@@
$\mathbb{E}[\exp(X^{t+1})]\leq\mathbb{E}\bigg[\hskip-1.99997pt\exp\hskip-1.99997pt\bigg(\hskip-3.00003pt\Big(1-\frac{a}{t+t_{0}}\Big)X^{t}+\hskip-3.99994pt\sum_{i\in[M]}\hskip-1.99997pt\frac{C_{i}}{(t+t_{0})^{i}}\hskip-1.99997pt\bigg)\hskip-1.99997pt\bigg].$
If $a\in(1,2]$ and $t_{0}\geq a$, we then have
@@
Lemma 5 provides a bound on the MGF of an “almost decreasing” process. Compared to bounds used in the centralized setting, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate, which is crucial for establishing linear speed-up, see Section 4 and the Appendix for detailed discussions. We are now ready to state the main result for strongly convex costs.
@@
Theorem 2 establishes HP convergence guarantees for smooth strongly convex costs and time-varying step-size, without requiring the bounded heterogeneity condition (A5). In the Appendix we provide the full bound, containing additional terms, of order $\mathcal{O}\big(\frac{\log(T+t_{0}))}{(T+t_{0})^{3}}\big)$. Next, we discuss multiple aspects of our results.
Centralized SGD and linear speed-up. Compared to centralized SGD, which achieves the HP rate $\mathcal{O}\Big(\frac{1}{T}+\frac{1}{T^{2}}\Big)$, e.g., [liu2024revisiting], the rate in Theorem 2 also consists of terms of different order, namely $\mathcal{O}\Big(\frac{1}{nT}+\frac{1}{T^{2}}\Big)$. Similarly to non-convex results, the leading term shows a linear speed-up in the number of users, again matching the corresponding MSE guarantees, e.g., [pmlr-v119-koloskova20a].
Network effect and transient time. We can see that network connectivity again only affects higher-order terms. Next, we can use the rate in Theorem 2 to establish a transient time of DSGD in the HP sense. As shown in the Appendix, our HP results imply a transient time of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$, which is strictly sharper than the transient time of DSGD in, e.g., [unified-refined, Table II], stemming from MSE rates, which is of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$. The sharper transient time further highlights the tightness of our results and an improved analysis.
Effects of noise and heterogeneity. While heterogeneity again affects only higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see the Appendix for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule. Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size. Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
Comparison with [dsgd-high-prob]. The authors in [dsgd-high-prob] establish HP convergence of DSGD for the more general case of PŁ costs, providing a rate of $\mathcal{O}\big(\frac{1}{T}\big)$, using the same time-varying step-size schedule considered in Theorem 2. Crucially, the authors require two very strong assumptions: $(i)$ path-wise uniformly bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq G$, for all $t\geq 1$ and some $G>0$, almost surely; $(ii)$ asymptotically vanishing noise, i.e., for all $i\in[n]$ and any $t\geq 1$, the authors require the following condition
@@
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Both conditions are very difficult to satisfy in practice, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is in general not satisfied even for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^11] It can be easily verified via Markov’s inequality that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, it follows from Jensen’s inequality that $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. Crucially, the authors in [dsgd-high-prob] are unable to establish linear speed-up in the number of users, even though much stronger assumptions on the noise and cost are imposed. On the other hand, our work uses standard assumptions, while simultaneously showing optimal convergence rates with linear speed-up and improved transient times compared to MSE results. These improvements stem not only from a sharper analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity assumption (A5) and ensure linear speed-up is achieved for strongly convex costs (see the Appendix for a detailed discussion).
@@ 4 Proof Outlines and Discussion
In this section we provide some proof sketches, outline the main challenges, and discuss our results.
Proof sketch of Theorem 1. Using Lemma 1, rearranging and summing up the first $T$ terms, we get
$\sum_{t\in[T]}\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2}\sum_{t\in[T]}\frac{\alpha_{t}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (4)
@@
We can use Lemma 3 to control the last term on the RHS. To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [li2020high, liu2023high, armacki2025optimal], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded. The rest of the proof then relies on Lemma 2, some technical results, careful selection of the step-size and Hölder’s inequality, see the Appendix for details.
@@
Proof sketch of Theorem 2. Starting from Lemma 1, subtracting $f^{\star}$ from both sides, using properties of strongly convex functions, defining $F^{t}\coloneqq n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$ and $A_{t}\coloneqq(t+t_{0}+1)\alpha_{t}$, we get
$F^{t+1}$ $\leq(1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}F^{t}-A_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}A_{t}nL\|\overline{z}^{t}\|^{2}+\frac{A_{t}L^{2}}{2}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (5)
We utilize the above inequality, Lemma 4, and a careful analysis to show that the MGF of $F^{t+1}$ satisfies the condition outlined in Lemma 5, after which we can use Lemma 5 to bound the MGF of the optimality gap. To complete the proof, we use the inequality [^12] See the Appendix for details.
$f(x^{t}_{i})-f^{\star}\leq 2(f(\overline{x}^{t})-f^{\star})+\frac{L}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2},$
@@
and apply our results to show that both optimality and consensus gaps are small, with high probability.
Discussion. Results of Theorems 1 and 2 are strong, demonstrating that DSGD converges in the HP sense under the same conditions on the cost function as in the MSE sense, while retaining order-optimal rates and linear speed-up in the number of users. Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we make several contributions. First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm. To that end, we provide Lemma 3 for non-convex costs, removing the uniformly bounded gradient condition and using the bounded heterogeneity condition (A5) instead, as well as the improved Lemma 4 for strongly convex costs, which removes (A5) altogether. Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense and is of independent interest for HP studies of decentralized methods. To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF of an “almost decreasing” process compared to existing (centralized) works, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], facilitating higher-order terms in the final bound, and is of independent interest (see the Appendix for a detailed discussion on the importance of Lemma 5 for achieving linear speed-up for strongly convex costs). Finally, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the resulting process is “almost decreasing”, in the sense of Lemma 5.
@@
On the dimension dependence. We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [pmlr-v119-koloskova20a, dsgd-high-prob]. This dependence stems from the third result in Lemma 2, where the dimension figures in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4). In particular, to show the third property in Lemma 2, we use the fact that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-norm-sub-Gaussian, as a direct proof fails to yield the variance reduction benefit in the number of users. One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [jin2019short]; see also the Appendix for a further discussion on the dimension dependence.

### v3

@@ 3.3 Strongly Convex Costs
In this subsection we provide results for the last iterate of strongly convex costs. Recall that strong convexity implies a unique global minimizer, e.g., [55], denoted by $x^{\star}\in\mathbb{R}^{d}$, with $f^{\star}\coloneqq f(x^{\star})$, and let $\kappa\coloneqq\frac{L}{\mu}$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively. It is known that DSGD does not require a heterogeneity bound of the form in (A5) for strongly convex costs in the MSE sense, e.g., [24, 33]. To show this is also the case in the HP sense, we next provide a refined version of Lemma 3, which carefully leverages properties of strong convexity to bound the MGF of the consensus gap.
@@
Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$, we have
@@
where $S_{k},D_{k}>0$ are problem related constants.
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5), with $S_{k},D_{k}$ depending on the step-size and other problem parameters, see Appendix F for explicit forms of $S_{k},D_{k}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional, geometrically decaying term in the exponent on the RHS, however, we use same initialization, for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which facilitates sharper bounds and ensures linear speed-up in the number of users.
@@
Let $\{X^{t}\}_{t\in\mathbb{N}}$ be a sequence of random variables initialized by a deterministic $X^{1}>0$, such that, for some $M\in\mathbb{N}$, $a,t_{0}>0$ and every $t\geq 1$
$\mathbb{E}[\exp(X^{t+1})]\leq\mathbb{E}\bigg[\hskip-1.99997pt\exp\hskip-1.99997pt\bigg(\hskip-3.00003pt\Big(1-\frac{a}{t+t_{0}}\Big)X^{t}+\hskip-3.99994pt\sum_{i\in[M]}\hskip-1.99997pt\frac{C_{i}}{(t+t_{0})^{i}}\hskip-1.99997pt\bigg)\hskip-1.99997pt\bigg],$
where $C_{i}>0$, $i\in[M]$. If $a\in(1,2]$ and $t_{0}\geq a$, we have
@@
Lemma 5 provides a tight bound on the MGF of an “almost decreasing” process, with the optimality gap of strongly convex costs being an instance of such a process. Compared to bounds used in the centralized setting, e.g., [21, 5, 41], Lemma 5 is significantly sharper and incorporates higher-order terms, which are crucial for ensuring linear speed-up, see Section 4 and Appendix I for detailed discussions. We are now ready to state the main result for strongly convex costs.
@@
Let (A1)-(A4) and (A6) hold, let the step-size schedule be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If constants $a$ and $t_{0}$ are chosen as $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
@@
Theorem 2 establishes HP convergence for smooth strongly convex costs and time-varying step-size. In Appendix F we provide the full bound, with additional higher-order terms. Next, we discuss multiple aspects of our results.
Centralized SGD and linear speed-up.
Compared to centralized SGD, which achieves the HP rate $\mathcal{O}\Big(\frac{1}{T}+\frac{1}{T^{2}}\Big)$, e.g., [41], the rate in Theorem 2 also consists of two terms of different order, namely $\mathcal{O}\Big(\frac{1}{nT}+\frac{1}{T^{2}}\Big)$. Similarly to non-convex results, the leading term shows a linear speed-up in the number of users, again matching the corresponding MSE guarantees, e.g., [33].
Network effect and transient time.
We can again see that network connectivity only affects higher-order terms. Morevoer, the rate in Theorem 2 implies a transient time of DSGD in the HP sense, of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. This is strictly sharper than the transient time of DSGD implied by MSE rates, of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see Table II in [1], further highlighting the tightness of our results and improved analysis.
Effects of noise and heterogeneity.
While heterogeneity only affects higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see Appendix F for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule. Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size. Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
Comparison with [44].
The authors in [44] establish HP convergence of DSGD for the more general case of PŁ costs, providing a rate of $\mathcal{O}\big(\frac{1}{T}\big)$, using the same time-varying step-size schedule as in Theorem 2. Crucially, the authors require two very strong assumptions: $(i)$ path-wise uniformly bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq G$, for some $G>0$ and all $t\geq 1$, almost surely; $(ii)$ asymptotically vanishing noise, i.e., for all $i\in[n]$ and any $t\geq 1$, they require the following
@@
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Both conditions are difficult to satisfy, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is not satisfied for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that it vanishes at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^9] Markov’s inequality implies that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, Jensen’s inequality implies $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. Crucially, [44] are unable to establish linear speed-up, even though they impose much stronger assumptions on the noise and cost. On the other hand, we show optimal convergence rate with linear speed-up using standard assumptions, with sharper transient times compared to MSE results. These improvements stem not only from a tighter analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity condition (A5) and ensure linear speed-up is achieved, see Appendix I for a further discussion.
Numerical experiments.
In Appendix G we provide numerical experiments which validate the main findings of the paper, notably showing that (i) DSGD converges in HP under the same conditions on the cost as in the MSE sense and (ii) DSGD achieves linear speed-up in the HP sense. The reader is kindly referred to the Appendix for details.
@@ 4 Proof Outlines and Discussion
In this section we briefly outline proof sketches of the main results and provide further discuss on our results.
Proof sketch of Theorem 1.
Using Lemma 1, rearranging and summing up the first $T$ terms, we get
$\sum_{t\in[T]}\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$
$+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2}\sum_{t\in[T]}\frac{\alpha_{t}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (4)
@@
We use Lemma 3 to control the last term on the RHS, while to deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ and remove the need for bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [37, 39, 2], i.e., we subtract $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of the inner product is absorbed by the left-hand side. The rest of the proof relies on Lemma 2, some technical results and a careful selection of the step-size, see Appendix E for details.
Proof sketch of Theorem 2.
@@
Using Lemma 1, properties of strongly convex functions, and defining $F^{t}\coloneqq n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$ and $A_{t}\coloneqq(t+t_{0}+1)\alpha_{t}$, we get
$F^{t+1}$ $\leq(1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}F^{t}-A_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$
$+\alpha_{t}A_{t}nL\|\overline{z}^{t}\|^{2}+\frac{A_{t}L^{2}}{2}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (5)
@@
We use (4), Lemma 4, and a careful analysis to show that the MGF of $F^{t}$ satisfies the conditions of Lemma 5, which we can then apply. The proof is completed by using these results to show that both optimality and consensus gaps are small with high probability, see Appendix F for details.
Comparison with centralized works.
Compared to works studying HP convergence of centralized SGD, e.g., [21, 5, 39, 41], we make several contributions. First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm. To that end, we provide Lemma 3 for non-convex, as well as the improved Lemma 4 for strongly convex costs, which removes the bounded heterogeneity condition (A5). Next, we establish Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense and is of independent interest when studying HP guarantees in decentralized settings. Finally, we provide a novel result on the MGF of the optimality gap of strongly convex costs (more broadly, “almost decreasing” processes), in the form of Lemma 5. This result provides a tighter control on the MGF compared to state-of-the-art bounds in [21, 5, 41] and is, in addition to Lemma 2, paramount to ensuring linear speed-up is achieved for strongly convex costs. In particular, following the same approach as in [21, 5], one can show that the MGF of the quantity of interest is uniformly bounded, i.e., that for some $B>0$ and all $t\geq 1$ and $\nu\in(0,B]$
$\mathbb{E}[\exp(\nu F^{t})]\leq\exp\Big(\frac{\nu}{B}\Big),$ (6)
where $F^{t}=n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$. Using (6) and Markov’s inequality, setting $\nu=B$, one can show that
$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}\leq\frac{\log(\nicefrac{{1}}{{\delta}})+1}{nB(t+t_{0})}\bigg)\geq 1-\delta,$ (7)
losing the linear speed-up, as $B=\mathcal{O}(1/n)$, which we show in Appendix I. On the other hand, Lemma 5 provides a much sharper bound on the MGF, ensuring that linear speed-up is achieved. Similar observations hold with respect to [41], see Appendix I for details.
On the dimension dependence.
@@
Our bounds contain a factor of $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for MSE or HP bounds in [33, 44]. This stems from Lemma 2, where $d$ appears in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4). One way to relax this dependence to $\log(d)$ is to assume that the noise is norm-sub-Gaussian, which comes with the cost of a slightly stronger noise condition, see, e.g., Corollary 7 in [27]; see also Appendix K for a further discussion on the dimension dependence.